OpenStudy (anonymous):

Find f(1) if f(x)=2x³+x²-3x-1

jimthompson5910 (jim_thompson5910):

\[\Large f(x)=2x^3+x^2-3x-1\] \[\Large f(1)=2(1)^3+(1)^2-3(1)-1\] \[\Large f(1)=2(1)+1-3(1)-1\] \[\Large f(1)=2+1-3-1\] \[\Large f(1)=-1\]